Grade 11Algebra

if w is the complex cube root of unity then w^2=?Please give me answer in detail

Abuzar Sheikh
9 Years agoGrade 11

1 Answer

Ashish bedi
9 Years ago
If w is complex cube root of unity then.1+w+w²=0 this is identity where 1,w,w² is cube root of unityW²= -1-wAnd w= -1/2+i√3/2 So the value of W²=-1/2-i√3/2.